OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 − 6x + 1 square feet, what is the length of one side of the room?

OpenStudy (anonymous):

1a) (3x - 1)² 1b) 4(9b² - 6b + 2) 1c) (4x + 3)² 1d) (2a - 5)² All polynomials are perfect square except 1b. 2. x² - 10x + 25 ==> (x - 5)² [A] 3. (3x + 7)² [C] 4. If the area of the square room is 4x² + 28x + 49, then.. A = s² s² = 4x² + 28x + 49 s² = (2x + 7)² s = 2x + 7 [B] Let's see if you can solve the last few problems by yourself.

hero (hero):

I don't even wanna how how you created those little squared symbols

hero (hero):

First time I seen that on OS

OpenStudy (anonymous):

its with the equation

hero (hero):

That's not what I meant. I know it's part of the equation

OpenStudy (anonymous):

oh will on mac there is a downloadable app that has all the equation stuff

hero (hero):

Isn't that nice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and its free so its nice to have

hero (hero):

What's the download link?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

for windows or mac?

hero (hero):

window

OpenStudy (anonymous):

im not sure they have it for windows i will see

hero (hero):

What's the name of the program?

OpenStudy (anonymous):

equation syimpels app

OpenStudy (anonymous):

you find it?

hero (hero):

wait

hero (hero):

nope, couldn't find it. How 'bout a link?

OpenStudy (anonymous):

its the mac app store

OpenStudy (anonymous):

its so confusing im only a fifth grader

OpenStudy (anonymous):

what lol why you doing this type of math then llol

OpenStudy (anonymous):

cause they skipped me up two grades

OpenStudy (anonymous):

O_O oh wow that sucks

OpenStudy (anonymous):

i know right please help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but i also have homework so one more qustion

hero (hero):

Please, I was doing these at 5th grade